OpenStudy (anonymous):

h - 48 + 17h = 42 h=?

OpenStudy (anonymous):

as stated before, 17h and h can combine to what? and then you want to add 48 to both sides and divide by the coefficient of h

OpenStudy (anonymous):

h=1 though right

OpenStudy (anonymous):

Can you please show us how you got that answer?

OpenStudy (anonymous):

So 18h + 48= 42 / what?

OpenStudy (anonymous):

\[\large h - 48 + 17h = 42\] \[\large h + 17h = 48+42\] \[\large 18h = 90\] \[\large h = 90/18 => 5\]

OpenStudy (anonymous):

What you do to one side, you must also do to the other.

OpenStudy (anonymous):